OpenStudy (anonymous):

Need Help please, Medal + Fan: Which equation represents an exponential function that passes through the point (2, 80)? f(x) = 4(x)5 f(x) = 5(x)4 f(x) = 4(5)x f(x) = 5(4)x

OpenStudy (mathstudent55):

Test the point in each equation and see which one works.

OpenStudy (anonymous):

How? could you explain please

OpenStudy (mathstudent55):

Also, an exponential function has the form \(\large f(x) = ab^x\) This alone will allow you to eliminate two answers.

OpenStudy (mathstudent55):

An exponential function has x as an exponent. What can you conclude about the first two choices?

OpenStudy (anonymous):

A and B cannot be correct

OpenStudy (mathstudent55):

Good. Now let's look at C and D.

OpenStudy (anonymous):

Okay..

OpenStudy (mathstudent55):

Choice C Is this true? \(80 = 4 \cdot 5^2\) Choice D Or is this true? \(80 = 5 \cdot 4^2\)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

16x5=80

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

Correct.